b) A man who has a mass of 80 kg can run up a flight of stairs of 20 steps, in 5 seconds. Given that each step is 20 cm high, how much power does the man use to run up the flight of stairs?

1
Expert's answer
2022-01-16T13:12:36-0500

Given:

"m=80\:\rm kg"

"n=20"

"t=\rm 5\: s"

"h_0=0.2\:\rm m"


The power of a man

"P=\frac{W}{t}=\frac{mgh}{t}=\frac{mgnh_0}{t}"

"P=\rm\frac{80*9.8*20*0.2}{5}=627\: J"